# Anomaly detection with flow-based fast calorimeter simulators

### v1

@@ V Conclusion and Outlook
By reweighting the proportion of showers originating from decays at each fixed displacement, we constructed signal datasets corresponding to fixed particle lifetimes. We found that CaloFlow has discriminative power for most of the models we tested. In particular, CaloFlow achieves the best performance in the case with $m_{\chi}=5$ GeV and $\tau=0.01$ ns where the particle is less highly boosted and majority of the particles decay close to the center of the detector.
Finally, we compared to performance of our unsupervised CaloFlow anomaly detector against a supervised classifier. We found that a supervised classifier trained on signal showers from highly boosted $\chi$ particles performed significantly poorer on showers from more massive, less highly boosted particles compared to our unsupervised method. On more boosted signals, the supervised classifier is still highly effective. This highlights the complementarity of different approaches and reaffirms the need to have a diversity of methods in order to achieve broad sensitivity.
@@
The datasets used in this study can be found at [52] and the software to generate these datasets are located at https://github.com/Ian-Pang/CaloGAN/tree/anomaly_detection. The machine learning software is at https://github.com/Ian-Pang/AD_with_CF.

### v2

@@ V Conclusion and Outlook
By reweighting the proportion of showers originating from decays at each fixed displacement, we constructed signal datasets corresponding to fixed particle lifetimes. We found that CaloFlow has discriminative power for most of the models we tested. In particular, CaloFlow achieves the best performance in the case with $m_{\chi}=5$ GeV and $\tau=0.01$ ns where the particle is less highly boosted and the majority of the particles decay close to the center of the detector.
Finally, we compared the performance of our unsupervised CaloFlow anomaly detector against a supervised classifier. We found that a supervised classifier trained on signal showers from highly boosted $\chi$ particles performed significantly poorer on showers from more massive, less highly boosted particles compared to our unsupervised method. When trained on signal showers from more massive $\chi$ particles and applied to signal showers from less massive $\chi$ particles, the supervised classifier still performs well. This highlights the complementarity of different approaches and reaffirms the need to have a diversity of methods in order to achieve broad sensitivity.
@@
The datasets used in this study can be found at [54] and the software to generate these datasets are located at https://github.com/hep-lbdl/CaloGAN/tree/samplingEM. The machine learning software is at https://github.com/Ian-Pang/AD_with_CF.